OpenStudy (anonymous):

simplify the expression. (9x^2 y^4 z^8)^-1/2

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

help me please

OpenStudy (misty1212):

the negative exponent means take the reciprocal, (flip it)

OpenStudy (misty1212):

the one half means take the square root

OpenStudy (misty1212):

\[\sqrt{9x^2y^4z^8}=9xy^2z^4\] so final answer is the reciprocal of that i.e \[\huge \frac{1}{3xy^2z^4}\]

OpenStudy (misty1212):

oops typo on the second line, but answer is correct

OpenStudy (anonymous):

thanks will you help me some more

OpenStudy (misty1212):

if i can go ahead and post